OpenStudy (anonymous):

find f'(X) when f(x)=cotx/1+cscx

OpenStudy (anonymous):

if it was me, i would start with \[\frac{\cos(x)}{1+\sin(x)}\] and take the derivative of that one using the quotient rule

OpenStudy (anonymous):

oh okay that makes it simple, thank you :)